# Conditional gene genealogies given the population pedigree for a diploid Moran model with selfing

### v1

@@ 1.2 Population genetics of selfing under neutrality
Our goal in this work is to understand how selfing affects the sampling structure of gene genealogies under neutrality when the process of coalescence is conditioned on the pedigree of the population. We ask whether results conditional on the pedigree differ from those of previous theoretical treatments which have disregarded this feature of the population. For simplicity we use a diploid Moran model of reproduction (Moran, 1958, 1962; Linder, 2009; Coron and Le Jan, 2022), which we will describe in Section 2. Previous theoretical treatments, including those we review here, have assumed Wright-Fisher reproduction (Fisher, 1930; Wright, 1931).
@@
Based on the considerations of Section 1.1, it may be that selfing species are even less likely than outcrossing ones to meet the assumptions of standard models of population genetics. Both the Wright-Fisher model and the Moran model apply to a single well mixed population of constant in size $N$ without selection. The simplest application relevant to selfing is to diploid, monoecious, randomly mating organisms, in which case selfing happens with probability $1/N$. When $N$ can be assumed to be large, both models converge to the Wright-Fisher diffusion with its corresponding coalescent process (Kingman, 1982; Hudson, 1983a; Tajima, 1983; Möhle, 1999; Ewens, 2004) which exist in the limit $N\to\infty$ with time measured in units proportional to $N$ generations. This provides a common framework for studying more complicated populations, as deviations from this simple null model. Selfing is one such deviation, which has been incorporated into the model by assuming that the selfing rate $s$ is a fixed constant when the population size $N$ tends to infinity.
@@
Pollak (1987) took a forward-time approach to studying, among other things, the equilibrium probabilities of identity by descent of two gene copies in the same individual or in different individuals, when $s\gg N^{-1}$. Let us call these probabilities $\Phi_{1}$ and $\Phi_{2}$, respectively. For large $N$, Pollak (1987) found that the expression for $\Phi_{2}$ was identical to that for a randomly mating population, if $N_{e}=N/(1+F)=(2-s)N/2$ is used in place of $N$. As $s$ increases from 00 to $1$, $N_{e}$ decreases from $N$ to $N/2$, effectively taking the value of $\Phi_{2}$ from that for a diploid (randomly mating) population to that for a haploid population. Pollak (1987) identified this as a separation-of-time-scales result, stemming from the assumption $s\gg N^{-1}$, and suggested that the two-times-scales diffusion approximation of Ethier and Nagylaki (1980) could be applied to prove convergence to the Wright-Fisher diffusion with this value of $N_{e}$. The separation of time scales can be seen in the relationship Pollak (1987) found between $\Phi_{1}$ and $\Phi_{2}$, namely that $\Phi_{1}=F+(1-F)\Phi_{2}$, which may best be understood in terms of subsequent descriptions of the coalescent process.
@@
It was in this context that Möhle (1998) proved an important general result (Möhle, 1998, Lemma 1, Theorem 1) for Markov processes with two time scales. Owing to the duality relation between the coalescent and diffusion models (Möhle, 1999) this result can be seen as the backward-time counterpart to the result of Ethier and Nagylaki (1980) cited by Pollak (1987). For the selfing model with fixed $s$ as $N\to\infty$, this gave rigorous justification to the model of Nordborg and Donnelly (1997): whenever a pair of lineages is in the same individual, they coalesce with probability $F=s/(2-s)$ instantaneously, otherwise ending up in different individuals with probability $1-F$, while samples from different individuals obey the standard neutral coalescent process with time in units of $(2-s)N$ generations (Möhle, 1998).
A somewhat different approach to the coalescent process with partial selfing was taken by Fu (1997), who approximated the single-generation backward-time transition probabilities for the number of ancestral lineages and the number of individuals containing them, up to order $1/N$, then used these in an algorithm to simulate coalescent ancestries. Time remains in generations, and the algorithm generates coalescence times of some few generations for pairs of gene copies within individuals. Something like this could be done by developing an algorithm for the pre-limiting ancestral process described in sections 3.2 and 3.3 in Möhle (1998). But it would not produce identical results, because Möhle (1998), also Nordborg and Donnelly (1997), used $s$ for the total proportion of selfing, while Fu (1997) used $s$ for the excess proportion of selfing, for a total of $s+(1-s)/N$.
1.3 Accounting for the pedigree of the population
The pedigree of a population is a graph representing the set of reproductive relationships of all individuals in the population for all time. Previous studies of the population-genetic effects of selfing, including those reviewed in Section 1.2, have followed standard practice in population genetics by averaging over all possible outcomes of reproduction, which is equivalent to averaging over the pedigree. See Diamantidis et al. (2024) for a recent detailed discussion of the issues surrounding this tradition in population genetics, which we review briefly here.
Averaging over outcomes of reproduction is often done implicitly, for example by stating that two ancestral lines coalesce with probability $1/(2N)$ each generation under the diploid monoecious Wright-Fisher model. Sometimes it is explicit, for example in equation (4.1) in Kingman (1982) for the coalescence probability under Cannings’ haploid exchangeable model (Cannings, 1974). For diploids in particular, the practice of averaging is highly questionable because all reproductive outcomes in the population for all time are encoded in a fixed structure, the genealogy or pedigree of the population through which all genetic transmission must also have occurred (Ball et al., 1990). In the most common application of coalescent models, which is to describe the distribution of gene genealogies or genetic variation among loci, the practice of averaging over reproductive outcomes is plainly wrong because the same pedigree holds for all loci.
A better-justified starting point for coalescent theory is to condition the ancestral genetic process on the pedigree. This gives the correct sampling structure for multiple unlinked loci: conditionally independent realizations of the process of genetic transmission within the same population/pedigree. The sampling structure implicit in traditionally formulated coalescent models is that each locus has its own pedigree, as if each locus was sampled from an independent population.
It turns out that the standard neutral coalescent model, at least, can still be applied because the coalescent model conditional on the pedigree is the same as the standard coalescent model for populations in the domain of the Kingman coalescent (Tyukin, 2015; Diamantidis et al., 2024). The reason is that pedigrees in large well mixed populations have a characteristic mixing time of $\log_{2}N$ generations (Chang, 1999; Derrida et al., 1999; Barton and Etheridge, 2011; Wakeley et al., 2012). But it is not true in general that extensions of the standard neutral coalescent model are similarly applicable to coalescent processes conditional on pedigrees generated under the corresponding population models (Wakeley et al., 2016; Wilton et al., 2017; Diamantidis et al., 2024).
@@ 1.3 Accounting for the pedigree of the population
The coalescent models of partial selfing in Fu (1997), Nordborg and Donnelly (1997) and Möhle (1998) average over the process of reproduction and thus implicitly over the pedigree. One consequence of instead conditioning on the pedigree is that each sampled individual has its own realized number of generations of selfing before the first outcrossing event in its ancestry. This, as we will show, is a single draw from the distribution (2). The end result of averaging is clear in previous approaches: $F$ in (3) is such an average. The averaging procedure is illustrated by the matrix $A^{m}$ in Möhle (1998, p. 496) which gives the $m$-generation transition probabilities of the fast process. For example,
@@
is the probability that the ancestors of two gene copies currently in the same individual remain distinct and are in two different individuals in past generation $m$. This can be expressed as a sum over the distribution (2), namely
@@
Partial selfing is just one of three possibilities when the coalescent process is conditioned on the pedigree. Here we reintroduce our notation, in which $\alpha_{N}$ denotes the selfing rate in a population of size $N$, and $\alpha$ denotes its value in the limit $N\to\infty$ or in the special case that $\alpha_{N}=\alpha$ is constant. When selfing is the primary mode of reproduction, an entirely new sort of model emerges in the limit, one which is completely invisible in the unconditional pedigree-averaging process. This novel limit process occurs when the outcrossing rate $1-\alpha_{N}$ is of order $N^{-1}$. It is parameterized by $\lambda\coloneq\lim_{N\to\infty}N(1-\alpha_{N})$, which we note is the same way mutation rates are assumed to scale in the Wright-Fisher diffusion and corresponding standard neutral coalescent model (Kingman, 1982; Hudson, 1983a; Tajima, 1983; Ewens, 2004) and how migration rates are scaled in the structured coalescent model (Takahata, 1988; Notohara, 1990; Herbots, 1997). In the limit we consider, $\lambda$ is the instantaneous rate of outcrossing events per ancestral lineage.
@@
Figure 1 shows simulation results of pedigrees and pairwise coalescence times for the model of selfing used by Möhle (1998), which corresponds to $\alpha_{N}=\alpha$ constant in our model. Specifically, for each of three selfing rates, $50$ independent pedigrees were simulated and for each of these two different individuals were sampled at random. Colored lines, one for each pedigree-plus-sample, depict the cumulative distribution function (CDF) of the coalescence time for two gene copies, one from each of the two sampled individuals. On the scale used in Figure 1, the corresponding expectation for the pedigree-averaging model (not depicted) would be the CDF of an exponential distribution with parameter $(2-\alpha)^{-1}$. Many of the CDFs in Figure 1(a) have something close to this shape, but those in Figure 1(b) and Figure 1(c) look nothing like it. Instead they are characterized by random jumps in probability at random times.
   Refer to caption   Refer to caption
@@
Figure 1: Cumulative distribution functions of the coalescence times for a sample of size $n=2$ on each of $50$ pedigrees simulated under the Wright-Fisher model with partial selfing and $N=1000$ individuals. Panels (a), (b) and (c) show the results for three different values of the probability of selfing, respectively, for $\alpha\in\{0.9,0.99,0.999\}$. Coalescence probabilities in each generation given the pedigree and the sampled individuals were calculated numerically by the method described in Wakeley et al (2012).
@@
The deviations from the predictions of the unconditional coalescent model displayed in Figure 1 are not due to the relatively small population size used in these simulations ($N=1000$). Instead they illustrate that the behavior we will prove for our Moran model occurs when $N(1-\alpha_{N})$ remains finite in the limit $N\to\infty$ is relevant even for such small populations. The value of $\alpha=0.99$ used in Figure 1(b) is similar to estimates of the selfing rates for the model species Arabidopsis thaliana and Caenorhabditis elegans (Abbott and Gomes, 1989; Sellinger et al., 2020; Barrière and Félix, 2005). Of course, $N=1000$ would be a very small population size for either of these species, though it may be realistic for some local populations. In any case, species or populations for which $N(1-\alpha)$ is not very large will have genetic ancestries very unlike what is expected under the unconditional or pedigree-averaging model. Instead, they are captured by the conditional coalescent model with ‘limited outcrossing’ which is characterized by a system of coalescing random walks on a random ancestral graph, described in Section 4.
@@ 1.4 Plan of the present work
Our analysis is based on a diploid Moran model of reproduction with constant population size $N$ and selfing rate $\alpha_{N}$, described in detail in Section 2. We focus on the coalescence time for a sample of size $n=2$, which already reveals the non-trivial effects of the pedigree on gene genealogies.
To establish the conditional limit for the limiting-outcrossing regime, we prove that the trajectories of potential ancestral individuals of the sample converge weakly to the ancestral graph, which is different from the proof for the partial-selfing regime. For the latter we adapt the method of Diamantidis et al. (2024), which itself was adapted from Birkner et al. (2013). We convert the problem of conditional convergence to $L^{2}$ convergence for the distribution of the minimum of two conditionally independent pairwise coalescence times for two independently assorting loci on the pedigree. Unlike the work in Diamantidis et al. (2024), here in both the conditional and unconditional case Möhle’s result (Möhle, 1998, Lemma 1, Theorem 1) does not apply, at least not in the obvious way. Some of our results could be obtained using the extension by Möhle and Notohara (2016), but here we take an alternate approach. We compute coalescence times by examining the random number of opportunities for coalescence between genetic lineages in different individuals, which we will refer to as overlap events, and the timings of these in relation to splitting events in the ancestral process, which result from outcrossing.
The organization of the paper is as follows. In Section 2, after detailing our Moran model, we make precise the pedigree and the gene genealogy. Our results for the unconditional and the conditional distributions of pairwise coalescence times are presented in Section 3 and Section 4, respectively. While we focus on $n=2$ in this paper, we briefly discuss the case $n>2$ and offer a conjecture in Section 4.2. In Section 5, we establish some fundamental properties of the scaling limit under limited outcrossing. This scaling limit is a coalescing random walk on a random ancestral graph. Our main convergence theorems will be proved in Section 6.
@@ 2 A discrete-time diploid Moran model with selfing
We consider a diploid, monoecious, panmictic (well mixed and randomly mating) population of constant $N$ over time. Individuals are labeled $I=\{1,...,N\}$. The population has overlapping generations and changes at discrete time-steps. At each time-step, a randomly chosen individual dies and is replaced by an offspring produced in one of two ways. With probability $1-\alpha_{N}$ the offspring has two distinct parents; with probability $\alpha_{N}$ the offspring has exactly one parent. The parent(s) are chosen uniformly at random from the previous time-step. In the case of two parents, these are chosen without replacement. The reproduction events at different time-steps are independent and identically distributed.
@@
    (distinct parents) With probability $1-\alpha_{N}$, a single offspring is produced by outcrossing. A triplet containing the two parents and one offspring $(\pi_{k,1},\,\pi_{k,2},\,\kappa_{k})\in I^{3}$ is chosen uniformly at random with $\pi_{k,1}\neq\pi_{k,2}$ (i.e. the parents are distinct). The offspring $\kappa_{k}$ is an individual in time-step $k$, while the offspring’s parents are the individuals in time-step $k+1$ with labels $\pi_{k,1}$ and $\pi_{k,2}$.
@@
    ($\kappa_{k}=3$)   time-step k($\pi_{k,1}=2,\pi_{k,2}=4$)  time-step k+1
@@
    (selfing) With probability $\alpha_{N}$, a parent-offspring pair with labels $(\pi_{k},\,\kappa_{k})\in I^{2}$ is chosen uniformly at random without replacement. The offspring $\kappa_{k}$ in time-step $k$ is the offspring of a single individual (the parent) with label $\pi_{k}$ at time-step $k+1$. This can be viewed as having $\pi_{k,2}=\pi_{k,1}$ in the previous reproduction event.
@@
    ($\kappa_{k}=3$)    time-step k($\pi_{k}=4$)   time-step k+1
The reproduction events at different time-steps are independent. Note that at time-step $k$, the offspring $\kappa_{k}$ is a new individual introduced and it replaces the one with the same label in the previous step $k+1$ (death-birth for the label $\kappa_{k}$); all the remaining $N-1$ individuals in time-step $k$ are the same individuals as those in time-step $k+1$. Hence this model has overlapping generations.
@@
When $\alpha_{N}=0$, our model is exactly the one introduced in Coron and Le Jan (2022). When $\alpha_{N}$ is equal to a fixed $p$ in $[0,1)$, our model is exactly that in Linder (2009).

### v2

@@ 1.2 Standard neutral models with a fixed selfing probability but no pedigree
Our goal in this work is to understand how selfing affects the sampling structure of gene genealogies under neutrality when the process of coalescence is conditioned on the pedigree of the population. For simplicity we use a diploid Moran model of reproduction (Moran, 1958, 1962; Linder, 2009; Coron and Le Jan, 2022), which we describe in Section 2. Here we review some relevant previous theoretical treatments which do not include pedigrees and which assume Wright-Fisher reproduction (Fisher, 1930; Wright, 1931).
@@
Both the Wright-Fisher model and the Moran model assume a single well mixed population of constant in size $N$ without selection. In the usual simplest application, i.e. to diploid, monoecious, randomly mating organisms, selfing occurs with probability $1/N$. When $N$ can be assumed to be large, both converge to the Wright-Fisher diffusion with its corresponding coalescent process (Ewens, 2004), which exist in the limit $N\to\infty$ with time measured in units proportional to $N$ generations. Arbitrary levels of selfing have been incorporated by assuming a fixed selfing probability of $s\in[0,1]$ in the limit $N\to\infty$. Based on the considerations of Section 1.1, it may be that selfing species are even less likely than outcrossing ones to meet the assumptions of these standard models of population genetics, but they provide a starting point for studying more complicated populations.
@@
Pollak (1987) took a forward-time approach to studying, among other things, the equilibrium probabilities of identity by descent of two gene copies in the same individual or in different individuals, when $s\gg N^{-1}$. Let us call these probabilities $\Phi_{1}$ and $\Phi_{2}$, respectively. For large $N$, Pollak (1987) found that the expression for $\Phi_{2}$ was identical to that for a randomly mating population, if $N_{e}=N/(1+F)=(2-s)N/2$ is used in place of $N$. As $s$ increases from 00 to $1$, $N_{e}$ decreases from $N$ to $N/2$, effectively taking the value of $\Phi_{2}$ from that for a diploid (randomly mating) population to that for a haploid population. Pollak (1987) identified this as a separation-of-time-scales result, stemming from the assumption $s\gg N^{-1}$, and suggested that the two-times-scales diffusion approximation of Ethier and Nagylaki (1980) could be applied to prove convergence to the Wright-Fisher diffusion with this value of $N_{e}$. The separation of time scales can be seen in the relationship Pollak (1987) found between $\Phi_{1}$ and $\Phi_{2}$, namely that $\Phi_{1}=F+(1-F)\Phi_{2}$, which may best be understood in terms of subsequent descriptions of the following coalescent process.
@@
It was in this context of partial selfing that Möhle (1998) proved an important general result (Möhle, 1998, Lemma 1, Theorem 1) for Markov processes with two time scales. Owing to the duality relation between the coalescent and diffusion models (Möhle, 1999) this result can be seen as the backward-time counterpart to the result of Ethier and Nagylaki (1980) cited by Pollak (1987). For the selfing model with fixed $s$ as $N\to\infty$, this gave rigorous justification to the model of Nordborg and Donnelly (1997): whenever a pair of lineages is in the same individual, they coalesce with probability $F=s/(2-s)$ instantaneously, otherwise ending up in different individuals with probability $1-F$, while samples from different individuals obey the standard neutral coalescent process with time in units of $(2-s)N$ generations (Möhle, 1998).
1.3 Accounting for the pedigree of the population is important
@@
The coalescent models of partial selfing in Fu (1997), Nordborg and Donnelly (1997) and Möhle (1998) average over the process of reproduction and thus implicitly over the pedigree. Clearly, one consequence of conditioning on the pedigree should be that each sampled individual has its own realized number of generations of selfing before the first outcrossing event in its ancestry. This, as we will show, is a single draw from the distribution (2). Previous approaches instead used the average $F$ in (3). The averaging procedure is illustrated by the matrix $A^{m}$ in Möhle (1998, p. 496) which gives the $m$-generation transition probabilities of the fast process. For example,
@@
is the probability that the ancestors of two gene copies currently in the same individual remain distinct and are in two different individuals in past generation $m$. This can be expressed as a sum over pedigrees using the distribution (2), namely
@@
Partial selfing is just one of three possibilities when the coalescent process is conditioned on the pedigree. Here we reintroduce our notation, in which $\alpha_{N}$ denotes the selfing probability in a population of size $N$, and $\alpha$ denotes its value in the limit $N\to\infty$. When selfing is the primary mode of reproduction, an entirely new sort of model emerges in the limit, one which is completely invisible in the unconditional pedigree-averaged process. This novel limit process occurs when the outcrossing probability $1-\alpha_{N}$ is of order $1/N$. It is parameterized by $\lambda\coloneq\lim_{N\to\infty}N(1-\alpha_{N})$, which we note is the same way mutation rates are assumed to scale in the Wright-Fisher diffusion and corresponding standard neutral coalescent model (Ewens, 2004) and how migration rates are scaled in the structured coalescent model (Takahata, 1988; Notohara, 1990; Herbots, 1997). In the limit we consider, $\lambda$ is the instantaneous rate of outcrossing events per ancestral lineage.
@@
Figure 1 shows simulation results of pedigrees and pairwise coalescence times for the model of selfing used by Möhle (1998), which corresponds to $\alpha_{N}=\alpha$ constant in our model. Specifically, for each of three selfing probabilities, $50$ independent pedigrees were simulated and for each of these two different individuals were sampled at random. Colored lines, one for each pedigree-plus-sample, depict the cumulative distribution function (CDF) of the coalescence time for two gene copies, one from each of the two sampled individuals. On the scale used in Figure 1, the corresponding expectation for the pedigree-averaging model in all three cases would be the CDF of an exponential distribution with parameter $(2-\alpha)^{-1}$. Many of the CDFs in Figure 1(a) have something close to this shape, but those in Figure 1(b) and Figure 1(c) look nothing like it. Instead they are characterized by random jumps in probability at random times. For any one population, these times are fixed and determine the predicted distribution of coalescence times among unlinked loci.
   Refer to caption   Refer to caption by the method in Wakeley et al. (2012).
@@
Figure 1: Cumulative distribution functions of the coalescence times for a sample of size $n=2$ on each of $50$ population pedigrees simulated under the Wright-Fisher model with partial selfing and $N=1000$ individuals. Panels (a), (b) and (c) show the results for three different values of the probability of selfing, respectively, for $\alpha\in\{0.9,0.99,0.999\}$. Coalescence probabilities in each generation given the pedigree and the sampled individuals were calculated exactly, up to numerical precision
@@
The deviations from the predictions of the unconditional coalescent model displayed in Figure 1 are not due to the relatively small population size used in these simulations ($N=1000$). Instead they illustrate that what we will prove is true in the limit $N\to\infty$ with $N(1-\alpha_{N})\to\lambda$ holds even for such small populations. The value $\alpha=0.99$ in Figure 1(b) is similar to estimates of the selfing probabilities for the model species Arabidopsis thaliana and Caenorhabditis elegans (Abbott and Gomes, 1989; Sellinger et al., 2020; Barrière and Félix, 2005). Of course, $N=1000$ would be a very small population size for either of these species, though it may be realistic for some local populations. In any case, species or populations for which $N(1-\alpha)$ is not very large will have genetic ancestries very unlike what is expected under the unconditional or pedigree-averaging model. Instead, they are captured by the conditional coalescent model with ‘limited outcrossing’ described in Section 4.
@@ 1.4 Plan of the present work
Our analysis is based on a diploid Moran model of reproduction with constant population size $N$ and selfing probability $\alpha_{N}$, described in detail in Section 2. We focus on the coalescence time for a sample of size $n=2$ at a single locus, which already reveals the non-trivial effects of the pedigree on gene genealogies. In the case of partial selfing, we prove convergence in distribution of the coalescence time via an analysis of two independently assorting or unlinked loci on the population pedigree, cf. Birkner et al. (2013) and Diamantidis et al. (2024). In the case of limited outcrossing, we take a different approach to establish the conditional limit. Here we focus on the set of diploid individual ancestors of the sample and prove that this converges weakly to an ancestral graph.
The organization of the paper is as follows. In Section 2, after detailing our Moran model, we make precise the pedigree and the gene genealogy. Our results for the unconditional and the conditional distributions of pairwise coalescence times are presented in Section 3 and Section 4, respectively. We briefly discuss the case $n>2$ and offer a conjecture in Section 4.1. In Section 5, we establish some fundamental properties of the scaling limit under limited outcrossing and of the associated ancestral graph. We present the proofs of our results in the Appendix, Section 7.
@@ 2 A discrete-time diploid Moran model with selfing
We consider a diploid, monoecious, panmictic (well mixed and randomly mating) population of constant size $N$. Individuals are labeled $I=\{1,...,N\}$. The population has overlapping generations and changes at discrete time-steps. At each time-step, a randomly chosen individual dies and is replaced by an offspring produced in one of two ways. With probability $1-\alpha_{N}$ the offspring has two distinct parents; with probability $\alpha_{N}$ the offspring has exactly one parent. The parent(s) are chosen uniformly at random from the previous time-step. In the case of two parents, these are chosen without replacement. The reproduction events at different time-steps are independent and identically distributed.
@@
    (distinct parents) With probability $1-\alpha_{N}$, a single offspring is produced by outcrossing. A triplet containing the two parents and one offspring $(\pi_{k,1},\,\pi_{k,2},\,\varepsilon_{k})\in I^{3}$ is chosen uniformly at random with $\pi_{k,1}\neq\pi_{k,2}$ (i.e. the parents are distinct). The offspring $\varepsilon_{k}$ is an individual in time-step $k$, while the offspring’s parents are the individuals in time-step $k+1$ with labels $\pi_{k,1}$ and $\pi_{k,2}$.
@@
    ($\varepsilon_{k}=3$)   time-step $k$($\pi_{k,1}=2,\pi_{k,2}=4$)  time-step $k$$+$$1$
@@
    (selfing) With probability $\alpha_{N}$, a parent-offspring pair with labels $(\pi_{k},\,\varepsilon_{k})\in I^{2}$ is chosen uniformly at random without replacement. The offspring $\varepsilon_{k}$ in time-step $k$ is the offspring of a single individual (the parent) with label $\pi_{k}$ at time-step $k+1$.
@@
    ($\varepsilon_{k}=3$)    time-step $k$($\pi_{k,1}=\pi_{k,2}=4$)   time-step $k$$+$$1$
Reproduction events in different time-steps are independent. Note that at time-step $k$, the offspring $\varepsilon_{k}$ is a new individual introduced and it replaces the one with the same label in the previous step $k+1$ (death-birth for the label $\varepsilon_{k}$). The remaining $N-1$ individuals in time-step $k$ are the same individuals as those in time-step $k+1$. Hence this model has overlapping generations.
@@
When $\alpha_{N}=0$, our model is exactly the one introduced in Coron and Le Jan (2022). When $\alpha_{N}$ is equal to a fixed $p$ in $[0,1)$, our model is exactly the one in Linder (2009).